OpenStudy (anonymous):

Find the next term of the following sequence. 25, 10, 4, ...

Parth (parthkohli):

Common ratio looks like 0.4. Multiply 0.4 to 4 to get the next term

Parth (parthkohli):

How did I get that common ratio? simple: Divide any term by the previous one

Parth (parthkohli):

10/25 = 0.4 That's it

OpenStudy (anonymous):

A.0 B.-5 C.8/5

ganeshie8 (ganeshie8):

ok.. here common ratio is 10/25 = 2/5 = 0.4 multiply 4 with 2/5 to get the next term :)

OpenStudy (anonymous):

sooo its what?